40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 5 15:28:38 2014, Steve, Update, SUS, HeNe laser test preparation beforeRSMqpdMoved.jpgafterSRMqpdMoved.jpg1103Ptest.pngND1onQPD.png
    Reply  Fri Sep 5 16:25:36 2014, Steve, Update, SUS, center ITMY oplev 
    Reply  Tue Sep 9 08:22:28 2014, Steve, Update, SUS, HeNe laser test preparation HeNecompared.png
       Reply  Tue Sep 9 16:59:32 2014, Steve, Update, SUS, HeNe laser test preparation 1103PqpdTranss.png
          Reply  Wed Sep 10 14:58:58 2014, Jenne, Update, SUS, OpLev test: New channels 
             Reply  Thu Sep 11 17:12:42 2014, Steve, Update, SUS, OpLev test: old SP qpd connected 
                Reply  Fri Sep 12 16:30:48 2014, Steve, Update, SUS, HeNe laser test preparation ND2.pngqpd222ND21103P.jpg
                   Reply  Wed Sep 17 13:41:00 2014, Jenne, Update, SUS, New calibrated channels for test QPD NewQPDcalibratedChans.png
                      Reply  Thu Sep 18 09:17:47 2014, Steve, Update, SUS, HeNe laser test  calibrated_16hrs.pngQPD222cal.jpg1103P40s40min.png
                         Reply  Mon Sep 22 10:14:07 2014, Steve, Update, SUS, HeNe laser test  1103Pmon3d.png
                            Reply  Tue Sep 23 08:34:14 2014, Steve, Update, SUS, HeNe laser test  tempDrop.png
                               Reply  Thu Sep 25 10:53:32 2014, Steve, Update, SUS, HeNe laser test  centeredDay6.png
          Reply  Wed Sep 10 15:31:16 2014, Steve, Update, SUS, HeNe laser test preparation SRMqpdisBack0nSRMsus.pngslowVSfast.png
    Reply  Wed Oct 15 17:09:49 2014, manasa, Update, General, Diode laser test preparation 
       Reply  Wed Oct 29 18:07:28 2014, manasa, Update, General, Diode laser test preparation 
Message ID: 10488     Entry time: Wed Sep 10 14:58:58 2014     In reply to: 10479     Reply to this: 10496
Author: Jenne 
Type: Update 
Category: SUS 
Subject: OpLev test: New channels 

Steve and EricG are moving their oplev test for aLIGO over to the SP table, so that we can have the SRM optical lever back.  

I have pulled out an Ontrak PSM2-10 position sensor and accompanying driver for the sensor.  This, like the POP QPD, has BNC outputs that we can take straight to the ADC.

In the c1pem model I have created 3 new filter modules:  C1:PEM-OLTEST_X, C1:PEM-OLTEST_Y, and C1:PEM-OLTEST_SUM.  I built, installed and restarted the model, and also restarted the daqd process on the frame builder.  On the AA breakout board on the 1X7 rack, these correspond to:

BNC # 29 = OLTEST_X

BNC # 30 = OLTEST_Y

BNC # 31 = OLTEST_SUM

By putting 1Vpp, 0.1Hz into each of these channels one at a time, I see on StripTool that they correspond as I expect.

Everything should be plug-and-play at this point, as soon as Steve is ready with the hardware.

ELOG V3.1.3-